OpenStudy (anonymous):

Help please? Which equation represents the line passing through the point (4, −16) and is perpendicular to the line y equals negative two thirds x plus 8?

OpenStudy (amistre64):

where are you stuck at?

OpenStudy (anonymous):

How do I draw a line when I only have one point? There's probably a way but I'm not sure how

OpenStudy (amistre64):

you are given a point and a 'slope'. the slope is just hidden at the moment in the information

OpenStudy (anonymous):

Well perpendicular means they cross at a right angle right?

OpenStudy (amistre64):

perpendicular lines have the property that their slopes are negative reciprocals of one another: if slope1 = 4/5 then the perp slope is -5/4

OpenStudy (amistre64):

yes, they cross at right angles

OpenStudy (anonymous):

So I should graph the line they give me first I guess

OpenStudy (amistre64):

you are given a line with a slope of -2/3 and want to apply the perp slope to your solution. what is the perp slope of -2/3?

OpenStudy (anonymous):

3/2?

OpenStudy (amistre64):

yes now, the line they ask for has a slope of: 3/2 and passes thru the point: (4,-16) we have all the information now to make the line equation.

OpenStudy (anonymous):

y = 3/2x... how do I find out what comes after that

OpenStudy (amistre64):

well, since we have a point and a slope; we could very well use what is called the point slope form of a line and then work it to whatever format they have as an option: given a point (a,b), and a slope, m ^^^^ ^^^^ point slope form is just: y-b = m(x-a) which we can rework as: y = mx - ma + b if need be to get it into slope intercept form

OpenStudy (anonymous):

Well the answers are a. y = 2/3x - 22 b. y = -2/3x - 22 c. y = -3/2x - 22 d. y = 3/2x - 22

OpenStudy (amistre64):

then it looks like d is our most reasonable choice.

OpenStudy (anonymous):

I don't understand where they got the -22 from though

OpenStudy (amistre64):

-3(4)/2 - 16 = 22

OpenStudy (amistre64):

.. = -22

OpenStudy (anonymous):

Ooh I understand thank you!!

OpenStudy (amistre64):

\[y-(-16)=\frac 32(x-(4))\] \[y=\frac 32x-(4)\frac{3}{2}-16\] \[y=\frac 32x-6-16\] ...

OpenStudy (amistre64):

but since all the options end in -22, that part doesnt really matter to us, the 3/2 slope is the key to simply picking out the right option without all the hulla-ballou

OpenStudy (anonymous):

Yeah! thank you for your help

OpenStudy (anonymous):

option C....